# CheXmask: a large-scale dataset of anatomical segmentation masks for multi-center chest x-ray images

### v1

@@ Abstract
The development of successful artificial intelligence models for chest X-ray analysis relies on large, diverse datasets with high-quality annotations. While several databases of chest X-ray images have been released, most include disease diagnosis labels but lack detailed pixel-level anatomical segmentation labels. To address this gap, we introduce an extensive chest X-ray multi-center segmentation dataset with uniform and fine-grain anatomical annotations for images coming from six well-known publicly available databases: CANDID-PTX, ChestX-ray8, Chexpert, MIMIC-CXR-JPG, Padchest, and VinDr-CXR, resulting in 676,803 segmentation masks. Our methodology utilizes the HybridGNet model to ensure consistent and high-quality segmentations across all datasets. Rigorous validation, including expert physician evaluation and automatic quality control, was conducted to validate the resulting masks. Additionally, we provide individualized quality indices per mask and an overall quality estimation per dataset. This dataset serves as a valuable resource for the broader scientific community, streamlining the development and assessment of innovative methodologies in chest X-ray analysis. The CheXmask dataset is publicly available at: https://physionet.org/content/chexmask-cxr-segmentation-data/.
@@
HybridGNet, a deep learning model for realistic organ contouring, offers a solution for the generation of anatomically plausible CXR segmentations [7, 8]. Utilizing a hybrid approach, it combines conventional convolution operations for image encoding with graph generative models for the anatomically-guided delineation of organ contours. The HybridGNet model was initially introduced with a small CXR landmark dataset to demonstrate its efficacy. In this work, we leverage this model to accomplish our main objective: introducing a large-scale segmentation dataset, named CheXmask, which provides anatomical masks with their corresponding quality index, for 6 extensive chest X-ray databases: CANDID-PTX[9], Chest x-ray8[10], Chexpert[2], MIMIC-CXR-JPG[11], Padchest[12] and VinDr-CXR[13]. These databases collectively represent a wide variety of geographical locations, patient demographics, and disease spectra, enabling the development of a broad, diverse segmentation dataset.
@@ Methods
In this study, we utilized six extensive CXR datasets: CANDID-PTX, ChestX-ray8, CheXpert, MIMIC-CXR-JPG, Padchest, and VinDr-CXR. Figure 1 provides an overview of the complete study, which was repeated six times, one for each dataset individually. The CANDID-PTX dataset [9], which encompasses 19,237 anonymized X-ray images, was collected at Dunedin Hospital in New Zealand between 2010 and 2020. The ChestX-ray8 dataset [10] contains 112,120 frontal-view X-ray images from 30,805 unique patients, annotated with text-mined fourteen disease image labels. The CheXpert dataset [2] comprises 224,316 chest radiographs collected from Stanford Hospital between October 2002 and July 2017. The MIMIC-CXR-JPG dataset [11] includes 377,110 CXR images from 227,827 imaging studies involving 65,379 patients; these were collected at the Beth Israel Deaconess Medical Center Emergency Department in the United States from 2011 to 2016. The Padchest dataset [12] is composed of 160,861 images from over 67,000 patients, collected from Hospital San Juan in Spain from 2009 to 2017. Lastly, the VinDr-CXR dataset [13] consists of 18,000 manually annotated images gathered from two primary hospitals in Vietnam.
Ethical statement
The datasets used in this study uphold rigorous ethical standards as reported in the original publications associated to every database. The CANDID-PTX dataset, sourced from New Zealand, underwent a comprehensive anonymization process wherein all identifiable details were systematically expunged. Access to this dataset required a two-step ethics training process, involving an online course and signing a Data Use Agreement. While no explicit information about the ethical or anonymization processes was found in the original papers for the ChestX-ray8 and CheXpert datasets, their use in this study was in compliance with established ethical standards. For the MIMIC-CXR-JPG dataset from the United States, anonymization was accomplished according to the Health Insurance Portability and Accountability Act of 1996 (HIPAA) Safe Harbor provisions. Accessing this dataset required the completion of a dedicated human subjects research training course. The Padchest dataset, gathered from Spain, received approval from the Hospital San Juan’s institutional research committee. This dataset was comprehensively anonymized and de-identified by the local health authorities. Lastly, the VinDr-CXR dataset from Vietnam ensured the privacy of subjects by eliminating or randomizing all personally identifiable information. A rigorous automated and manual verification process was implemented to ensure the thorough removal of all textual information containing identifiable data. Accessing this dataset also required the completion of a training course. To ensure the ethical use of all datasets, the authors of this study have fulfilled all necessary requirements, including completion of requisite ethics courses and signing relevant data use agreements.
@@
The CANDID-PTX, ChestX-ray8, and VinDr-CXR datasets contain only frontal images, so all images were included. We incorporated images from the Padchest dataset based on the “Projection” metadata field, including only those labeled as PA or AP, reducing the initial count of 160,861 images to 96,287. For the CheXpert dataset, images labeled as “Frontal” in the “Frontal/Lateral” metadata column were included, resulting in 187,825 out of the total 224,316 images. Lastly, images from the MIMIC-CXR-JPG dataset were included if the “ViewPosition” tag was listed as “PA” or “AP”, which resulted in 243.334 out of the initial 377,110 images.
@@
For compatibility with HybdriGNet input format, we preprocessed images to attain a uniform size suitable for generating pseudo-landmark contours. The standard size of the training data was 1024x1024 pixels, obtained by padding the images to square dimensions then resizing them to the required size, when needed.
Images from ChestX-ray8 dataset already satisfied the required dimension and format, thus requiring no pre-processing. The CANDID-PTX dataset met the required image dimension in DICOM format, so we converted images to PNG. The datasets of CheXpert, MIMIC-CXR-JPG, Padchest, and VinDr-CXR underwent a standard pre-processing pipeline, which included padding the images to a square shape and then resizing them to 1024x1024 pixels. Additionally, VinDr-CXR images needed extraction from DICOM files before the pre-processing. All these pre-processing procedures are reversible, which allows for the preservation of the original image shapes for releasing the generated annotations.
@@
In this work, we used the modified multi-centric training procedure [15] to retrain the HybridGNet model with the complete dataset used in prior studies (i.e., both the train and test splits of previous works). This dataset, which will be referred to as Chest-Xray-Landmark dataset, was released previously and is available on GitHubaaahttps://github.com/ngaggion/Chest-xray-landmark-dataset. The Chest-Xray-Landmark dataset incorporates 911 annotated images from the JSRT [16], Montgomery [17], Shenzhen [18] datasets, and a subset of the Padchest dataset [12]. All images contain lung annotations and 383 images also include heart segmentations. The training process incorporated various data augmentation techniques, such as random rotations, scaling, and color shifting, to enhance robustness and generalizability.
@@
Figure 3: Histogram showcasing the distribution of the RCA-estimated DSC for the complete CheXmask database. Example images with their landmark-based segmentations were drawn with lines to their corresponding histogram bin.
@@
The CheXmask dataset is structured as one file of comma-separated values (CVS) for each CXR dataset. The CSV content is explained in Table 1. We do not include images or metadata from the original datasets. We include an image ID in the first column of the CSV files, which has the same column name as the ID column of the respective dataset, and allows to match the rows in this CSV with the original dataset’s instances. We also include the pre-processed version of the masks, allowing to use all datasets in the same image resolutions (please also note that CANDID-PTX and ChestX-ray8 masks are already in the desired resolution). This database [21] is available at PhysioNet [22], in the following repository: https://physionet.org/content/chexmask-cxr-segmentation-data/.
@@
The technical validation of the CheXmask dataset was conducted through a three-tiered approach: (1) evaluating the quality of HybridGNet segmentations by measuring the DSC on a gold-standard subset revised by an expert physician; (2) validating the relevance of the RCA-estimated DSC as performance metric; (3) evaluating the quality of the masks for the six source datasets separately, by exploring their suitability as a training GT set for the segmentation task.
@@
We built a gold-standard set of masks (for images sampled from the six original datasets) labeled by an experienced physician to evaluate the segmentation quality of CheXmask lungs and heart masks. The physician, counting with more than five years of experience, performed a manual revision of the right lung, left lung and heart landmark-based segmentations. We used an open-source labeling platform called LabelStudio to do so. The landmarks predicted by HybridGNet for a subset of images were uploaded as predictions to LabelStudio, where the annotator corrected these predicted landmarks by moving them to their correct specific contour positions. This guarantees that the GT masks are based on the same collection of nodes than the evaluated masks, and reduces the time needed for annotation. Figure 4 shows the annotation interface. We include the source code to launch this interface in our repository, as an easy tool to perform further validations. We measured the DSC of HybridGNet’s masks comparing with the manually-corrected landmark-annotations, and report statistics on the DSC distribution for the six public datasets.
@@
The sample size of the subset used as gold-standard evaluation was chosen based on the work by El Jurdi and Colliot [23], which suggests sampling 100 images for a 4%-wide confidence interval in segmentation tasks. We sampled 30 images per dataset, obtaining a final gold-standard set of 180 images. We combined random sampling and histogram-based sampling to ensure the presence of high and low quality masks. For each dataset, ten images were sampled based on a 10-bins histogram of the RCA-estimated DSC of all the dataset’s images, taking one random image per bin. Another twenty images were then randomly selected from the whole dataset.
@@
Figure 4: Label Studio setup for manual landmark-based segmentation
@@
Table 2 summarizes the results of physician validation for lung and heart segmentation across the six datasets on their 30-samples subsets. We report DSC, Hausdorff Distance (HD), and Hausdorff 95% Distance (HD95) as evaluation metrics. HD and HD95 are distance error metrics, providing insights into the difference between the contours of the GT masks and the predicted masks.
@@
In terms of lung segmentation, the datasets generally exhibit high DSC values, ranging from 0.947 to 0.981. This indicates a strong overlap between the GT masks and the predicted lung masks. Among the datasets, CANDID-PTX exhibits the highest agreement, while CheXpert has the lowest mean DSC. CheXpert shows relatively higher HD and HD95 values compared to other datasets, suggesting a larger discrepancy between the contours.
For heart segmentation, the DSC values range from 0.950 to 0.980, indicating a strong overlap between the GT and predicted heart masks across all datasets. The HD and HD95 values for heart segmentation are generally lower compared to lung segmentation, indicating a smaller distance between the contours.
@@
Overall, the results indicate that the segmentation models perform well in capturing the lung and heart structures. CANDID-PTX consistently demonstrates the highest agreement for both lungs and heart, while CheXpert exhibits relatively lower agreement. These findings highlight the variations in segmentation performance across different datasets and provide valuable insights for further improving the accuracy of the segmentation algorithms.
@@
Figure 5 illustrates the MSE per landmark across the whole dataset. Each landmark is color-coded based on its respective MSE value, using a logarithmic scale for better visualization. The size of each landmark also corresponds to the magnitude of the MSE, with larger landmarks indicating higher errors. As can be seen in this figure, the bottom part of the lungs and the heart landmarks tend to be larger, denoting higher prediction errors. This aligns with the observations made by the expert physician responsible for the manual annotations, who noted that the boundaries of the heart and the lower section of the lungs tend to exhibit more blurriness compared to the upper part.
Figure 5: Illustration of the Mean Squared Error (MSE) per landmark across the entire gold-standard set, depicted in logarithmic scale for improved visual clarity. The color intensity of each landmark represents the magnitude of the MSE. Moreover, the size of each landmark is proportional to its respective MSE, thus larger landmarks indicate a greater prediction error.
Table 2: Results of physician validation with standard deviations ($n=30$ per dataset)
Lungs Heart
Dataset DSC HD HD95 DSC HD HD95
CANDID-PTX 0.981 $\pm$ 0.038 23.7 $\pm$ 43.0 11.9 $\pm$ 23.6 0.950 $\pm$ 0.111 23.6 $\pm$ 51.9 21.0 $\pm$ 46.6
CheXpert 0.947 $\pm$ 0.087 39.3 $\pm$ 57.7 27.0 $\pm$ 46.3 0.961 $\pm$ 0.083 18.0 $\pm$ 37.2 14.8 $\pm$ 30.3
ChestX-Ray8 0.961 $\pm$ 0.097 32.3 $\pm$ 63.4 18.8 $\pm$ 44.1 0.958 $\pm$ 0.133 18.6 $\pm$ 58.6 17.2 $\pm$ 55.1
MIMIC-CXR-JPG 0.965 $\pm$ 0.077 29.7 $\pm$ 51.1 15.1 $\pm$ 32.2 0.954 $\pm$ 0.101 22.9 $\pm$ 51.9 21.2 $\pm$ 47.8
Padchest 0.966 $\pm$ 0.089 23.0 $\pm$ 42.7 12.7 $\pm$ 27.5 0.980 $\pm$ 0.035 12.8 $\pm$ 24.3 10.4 $\pm$ 18.8
VinDr-CXR 0.978 $\pm$ 0.048 18.8 $\pm$ 40.1 8.9 $\pm$ 18.5 0.953 $\pm$ 0.119 21.9 $\pm$ 55.4 20.6 $\pm$ 52.6
@@
Figure 7: Histograms illustrating the distribution of segmentation quality across the six datasets. Each histogram represents the distribution of RCA (Reverse Classification Accuracy) estimations of DSC for a specific dataset, providing a visual representation of the segmentation performance. The histograms are truncated at 0.65 for visualization purposes. The full range of values is considered in the statistical analysis (Table 3).
@@
Figure 8 presents histograms that clearly demonstrate a superior RCA-estimated DSC for PA images compared to AP images across all investigated datasets. This is consistent with the fact that PA images tend to come from hospitalized patients, who are more difficult to position in standard views and usually include artifacts or cables that may act as confounders or occlude body parts, making anatomical segmentation more challenging. In contrast, discrepancies based on sex were less pronounced, with a slightly diminished segmentation quality observed in females. This is also consistent with the fact x-ray imaging of the upper thorax women’s breasts tend to occlude the imaged organs, resulting in poorer image contrast for the relevant anatomy [26] and making the segmentation task more challenging. Furthermore, the influence of disease findings on segmentation quality varied amongst datasets, with significant disparities evident in the ChestX-ray8 and CheXpert datasets.
@@
Table 3 can be useful to estimate the prevalence of such out-of-distribution images in each dataset. For instance, in the CANDID-PTX and VinDr-CXR datasets, the number of out-of-distribution images is relatively low. Notably, the VinDr-CXR dataset does not include any non-posteroanterior (PA) images, as all were manually annotated by physicians. However, the inclusion of zoomed-out images capturing both arms, which were absent in the HybridGNet training set, leads to lower-quality segmentations in those cases. In contrast, the larger datasets contain a higher number of cases mislabeled as PA or AP in their metadata. These mislabeled images could be easily identified and discarded by thresholding the RCA-estimated DSC, for example using the 1% or 5% quartile of a dataset as threshold. This approach can help to improve the overall quality and reliability of downstream applications developed using this datasets.
Figure 9: Illustration of a random subset of images with some of the lowest RCA values from each dataset. These images were identified as having low-quality segmentations based on their corresponding RCA scores. The selection showcases various types of images, including those with noise, mislabeled lateral CXR, and even non-medical objects like a mobile phone in the Padchest dataset.
@@
We performed a final quality control at the dataset-level to evaluate the quality of HybridGNet segmentations per dataset (not individually as we did with RCA-estimated DSC). The hypothesis is that if the segmentations are good, they can be used as training set for a segmentation model that will obtain high performance. Thus, we trained six new HybridGNet models, by using the set of predicted masks of each of the six public datasets as training GT. We then evaluated model performance on the complete ChestXray-Landmarks database. All models were trained using the same configurations and hyperparameters as the original HybridGNet model [15]. As a comparison reference, we include the performance metrics reported on prior studies for the 20% test split of the ChestXray-Landmarks database which can be seen as an upper bound in performance, as this model was trained with the training split of the ChestXray-Landmarks database.
@@
The CheXmask dataset is intended to serve as a resource for researchers working in the field of medical imaging, particularly those interested in computational anatomy, shape understanding, CXR analysis and image segmentation. It may be useful for a wide range of downstream applications, including deep learning model development and evaluation, training of generative models, clinical decision support systems, disease detection and diagnosis, and anomaly detection.
For the sake of completeness, we release segmentation masks for all images included in our study. However, based on the qualitative and quantitative analyses presented in this paper (see section Technical Validation), our recommendation for downstream tasks is to only use segmentations whose RCA-estimated DSC is higher than 0.7, to avoid including out-of-distribution images as well as low quality masks.

### v3

@@ Abstract
The development of successful artificial intelligence models for chest X-ray analysis relies on large, diverse datasets with high-quality annotations. While several databases of chest X-ray images have been released, most include disease diagnosis labels but lack detailed pixel-level anatomical segmentation labels. To address this gap, we introduce an extensive chest X-ray multi-center segmentation dataset with uniform and fine-grain anatomical annotations for images coming from five well-known publicly available databases: ChestX-ray8, Chexpert, MIMIC-CXR-JPG, Padchest, and VinDr-CXR, resulting in 657,566 segmentation masks. Our methodology utilizes the HybridGNet model to ensure consistent and high-quality segmentations across all datasets. Rigorous validation, including expert physician evaluation and automatic quality control, was conducted to validate the resulting masks. Additionally, we provide individualized quality indices per mask and an overall quality estimation per dataset. This dataset serves as a valuable resource for the broader scientific community, streamlining the development and assessment of innovative methodologies in chest X-ray analysis. The CheXmask dataset is publicly available at: https://physionet.org/content/chexmask-cxr-segmentation-data/.
@@
HybridGNet, a deep learning model for realistic organ contouring, offers a solution for the generation of anatomically plausible CXR segmentations [7, 8]. Utilizing a hybrid approach, it combines conventional convolution operations for image encoding with graph generative models for the anatomically-guided delineation of organ contours. The HybridGNet model was initially introduced with a small CXR landmark dataset to demonstrate its efficacy. In this work, we leverage this model to accomplish our main objective: introducing a large-scale segmentation dataset, named CheXmask, which provides anatomical masks with their corresponding quality index, for 5 extensive chest X-ray databases: Chest x-ray8[9], Chexpert[2], MIMIC-CXR-JPG[10], Padchest[11] and VinDr-CXR[12]. These databases collectively represent a wide variety of geographical locations, patient demographics, and disease spectra, enabling the development of a broad, diverse segmentation dataset.
@@
In this context, it is important to notice that there are limited publicly available chest X-ray datasets with expert segmentation masks, and most studies mainly rely on the Shenzhen Hospital [14], Montgomery County [15], and JSRT [16] datasets. Taken together, they correspond to less than 1000 annotations and the segmented structures are not consistent (for example, while JSRT includes lung, hearth and clavicle annotations, Shenzhen and Montgomery only include lung annotations). Concurrently to our work, larger datasets have recently emerged, like the one released by Seibold and co-workers [17]. However, even though this dataset provides segmentations for a large number of X-ray images featuring different anatomical structures, they are not real X-ray images, since they were generated as digitally reconstructed radiographs, by projecting CT images from the PAX-Ray++ Dataset [18]. As highlighted by the authors, real X-ray images tend to use a harder type of radiation than standard CTs. This fact, together with the contrast agents used in CTs, might lead to a divergence of visual quality between the projected X-rays and real ones [17]. Moreover, they only provide pixel-level masks, while we also provide landmark based segmentations with correspondences. In this context, CheXmask stands out as a unique large scale dataset for real X-ray images, with quality control validated through objective indices.
@@ Methods
In this study, we utilized five extensive CXR datasets: ChestX-ray8 (available in a dedicated website [19]), CheXpert (available in a dedicated website [20]), MIMIC-CXR-JPG (available at PhysioNet[21]), Padchest (available in a dedicated website [22]), and VinDr-CXR (available at PhysioNet[23]), original sources to download each dataset are provided in the web references. Figure 1 provides an overview of the complete study, which was repeated five times, one for each dataset individually. The ChestX-ray8 dataset [9] contains 112,120 frontal-view X-ray images from 30,805 unique patients, annotated with text-mined fourteen disease image labels. The CheXpert dataset [2] comprises 224,316 chest radiographs collected from Stanford Hospital between October 2002 and July 2017. The MIMIC-CXR-JPG dataset [10] includes 377,110 CXR images from 227,827 imaging studies involving 65,379 patients; these were collected at the Beth Israel Deaconess Medical Center Emergency Department in the United States from 2011 to 2016. The Padchest dataset [11] is composed of 160,861 images from over 67,000 patients, collected from Hospital San Juan in Spain from 2009 to 2017. Lastly, the VinDr-CXR dataset [12] consists of 18,000 manually annotated images gathered from two primary hospitals in Vietnam.
@@
ChestX-ray8, and VinDr-CXR datasets contain only frontal images, so all images were included. We incorporated images from the Padchest dataset based on the “Projection” metadata field, including only those labeled as PA or AP, reducing the initial count of 160,861 images to 96,287. For the CheXpert dataset, images labeled as “Frontal” in the “Frontal/Lateral” metadata column were included, resulting in 187,825 out of the total 224,316 images. Lastly, images from the MIMIC-CXR-JPG dataset were included if the “ViewPosition” tag was listed as “PA” or “AP”, which resulted in 243.334 out of the initial 377,110 images.
@@
For compatibility with HybridGNet input format, we preprocessed images to attain a uniform size suitable for generating pseudo-landmark contours. The standard size of the training data was 1024x1024 pixels, obtained by padding the images to square dimensions then resizing them to the required size, when needed.
Images from ChestX-ray8 dataset already satisfied the required dimension and format, thus requiring no pre-processing. CheXpert, MIMIC-CXR-JPG, Padchest, and VinDr-CXR datasets underwent a standard pre-processing pipeline, which included padding the images to a square shape and then resizing them to 1024x1024 pixels. Additionally, VinDr-CXR images needed extraction from DICOM files before the pre-processing. All these pre-processing procedures are reversible, which allows for the preservation of the original image shapes for releasing the generated annotations.
@@
In this work, we used the modified multi-centric training procedure [24] to retrain the HybridGNet model with the complete dataset used in prior studies (i.e., both the train and test splits of previous works). This dataset, which will be referred to as Chest-Xray-Landmark dataset, was released previously and is available on GitHub: https://github.com/ngaggion/Chest-xray-landmark-dataset. The Chest-Xray-Landmark dataset is curated from a variety of sources, providing annotations for a total of 911 images. These images are sourced from different repositories, including the JSRT dataset [16] (available from its dedicated website [25]), a small subset of the Padchest dataset [11] (accessible on its designated website [22]), as well as the Montgomery [15] and Shenzhen [14] datasets (both obtainable from the NIH National Library of Medicine [26]). Please note that the Chest-Xray-Landmark dataset does not provide the images themselves; users are required to source the original images from their respective origins. All 911 images within this dataset contain lung annotations and only 383 images also include heart labels. The training process incorporated various data augmentation techniques, such as random rotations, scaling, and color shifting, to enhance robustness and generalizability.
A fully working demo of the model used to segment the complete dataset is available on Hugging Face Spaces at https://huggingface.co/spaces/ngaggion/Chest-x-ray-HybridGNet-Segmentation.
@@
Figure 3: Histogram showcasing the distribution of the RCA-estimated DSC for the complete CheXmask database. Example images with their landmark-based segmentations were drawn with lines to their corresponding histogram bin. Reproduction of x-ray images was allowed upon request to the original sources.
@@
The CheXmask dataset is structured as one file of comma-separated values (CVS) for each CXR dataset. The CSV content is explained in Table 1. We do not include images or metadata from the original datasets. We include an image ID in the first column of the CSV files, which has the same column name as the ID column of the respective dataset, and allows to match the rows in this CSV with the original dataset’s instances. We also include the pre-processed version of the masks, allowing to use all datasets in the same image resolutions (please also note that ChestX-ray8 masks are already in the desired resolution). This database [29] is available at PhysioNet [30], in the following repository: https://physionet.org/content/chexmask-cxr-segmentation-data/.
@@
The technical validation of the CheXmask dataset was conducted through a three-tiered approach: (1) evaluating the quality of HybridGNet segmentations by measuring the DSC on a gold-standard subset revised by an expert physician; (2) validating the relevance of the RCA-estimated DSC as performance metric; (3) evaluating the quality of the masks for the five source datasets separately, by exploring their suitability as a training GT set for the segmentation task.
@@
We built a gold-standard set of masks (for images sampled from the five original datasets) labeled by two experienced physicians, referred as P1 and P2, to evaluate the segmentation quality of CheXmask lungs and heart masks. Both physicians, counting with more than five years of experience, performed a manual revision of the right lung, left lung and heart landmark-based segmentations. We used an open-source labeling platform called LabelStudio to do so. The landmarks predicted by HybridGNet for a subset of images were uploaded as predictions to LabelStudio, where the annotator corrected these predicted landmarks by moving them to their correct specific contour positions. This guarantees that the GT masks are based on the same collection of nodes than the evaluated masks, and reduces the time needed for annotation. Figure 4 shows the annotation interface. We include the source code to launch this interface in our repository, as an easy tool to perform further validations. We measured the DSC of HybridGNet’s masks comparing with the manually-corrected landmark-annotations, and report statistics on the DSC distribution for the five public datasets.
@@
The sample size of the subset used as gold-standard evaluation was chosen based on the work by El Jurdi and Colliot [31], which suggests sampling 100 images for a 4%-wide confidence interval in segmentation tasks. We sampled 50 images per dataset, obtaining a final gold-standard set of 250 images. We combined random sampling, histogram-based sampling and a random selection of images which included radiological findings (disease or abnormalities) to ensure the presence of easy and challenging cases. For each dataset, ten images were sampled based on a 10-bins histogram of the RCA-estimated DSC of all the dataset’s images, taking one random image per bin. Another twenty images were then randomly selected from the whole dataset. Finally, 20 additional images were taken from each dataset, ensuring the presence of pathology or findings such as edema, consolidation, pelural effusion or support devices, which usually result in more challenging segmentation scenarios.
@@
Figure 4: Label Studio setup for manual landmark-based segmentation. Reproduction of x-ray images was allowed upon request to the original sources.
@@
Table 2 summarizes the results of physician validation for lung and heart segmentation across the five datasets on their 50-samples subsets. We report DSC, Hausdorff Distance (HD), and Hausdorff 95% Distance (HD95) as evaluation metrics. HD and HD95 are distance error metrics, providing insights into the difference between the contours of the GT masks and the predicted masks.
@@
In terms of lung segmentation, the datasets generally exhibit high DSC values, ranging from 0.949 to 0.982. This indicates a strong overlap between the GT masks and the predicted lung masks. Among the datasets, VinDr-CXR exhibits the highest agreement for both physicians, while CheXpert has the lowest mean DSC.
For heart segmentation, the DSC ranges differ between the different annotators, in one case from 0.957 to 0.979 and, in the other, from 0.912 to 0.948. In both cases there is a strong overlap between the GT and predicted heart masks across all datasets (heart segmentation is a more difficult task). The HD and HD95 values for heart segmentation are generally lower compared to lung segmentation, indicating a smaller distance between the contours. When comparing between the two expert physicians, the results vary in the same way as with the predictions. More difficult datasets such as CheXpert have the largest inter-reviewer differences, meanwhile VinDr-CXR has both better performance in each reviewer and the lowest inter-reviewer difference.
@@
Overall, the results indicate that the segmentation models perform well in capturing the lung and heart structures. VinDr-CXR consistently demonstrates the highest agreement for both lungs and heart, while CheXpert exhibits relatively lower agreement. These findings highlight the variations in segmentation performance across different datasets and provide valuable insights for further improving the accuracy of the segmentation algorithms.
@@
Figure 5 illustrates the MSE per landmark across the whole dataset per expert reviewer. Each landmark is color-coded based on its respective MSE value, using a logarithmic scale for better visualization. The size of each landmark also corresponds to the magnitude of the MSE, with larger landmarks indicating higher errors. As can be seen in this figure, the bottom part of the lungs and the heart landmarks tend to be larger, denoting higher prediction errors. This aligns with the observations made by both expert physicians responsible for the manual annotations, who noted that the boundaries of the heart and the lower section of the lungs tend to exhibit more blurriness compared to the upper part.
Figure 5: Illustration of the Mean Squared Error (MSE) per landmark across the entire gold-standard set per expert reviewer, depicted in logarithmic scale for improved visual clarity. The color intensity of each landmark represents the magnitude of the MSE. Moreover, the size of each landmark is proportional to its respective MSE, thus larger landmarks indicate a greater prediction error.
Table 2: Results of clinical validation. Comparison of predictions vs P1 and P2 ground truth annotations ($n=50$ per dataset)
DSC Lungs HD Lungs HD 95 Lungs DSC Heart HD Heart HD 95 Heart
Pred vs P1
CheXpert 0.949 (0.067) 49.9 (49.7) 30.0 (38.8) 0.912 (0.083) 40.8 (34.0) 36.4 (32.6)
ChestX-Ray8 0.966 (0.078) 35.4 (53.7) 18.6 (37.6) 0.921 (0.111) 34.8 (45.3) 30.9 (43.1)
MIMIC-CXR-JPG 0.962 (0.055) 37.5 (40.4) 21.1 (30.5) 0.919 (0.093) 36.1 (38.5) 32.5 (36.6)
Padchest 0.965 (0.067) 36.7 (35.9) 17.9 (23.0) 0.948 (0.048) 26.5 (23.1) 22.7 (20.6)
VinDr-CXR 0.982 (0.035) 27.2 (33.0) 10.2 (15.4) 0.912 (0.098) 39.8 (42.8) 35.8 (41.3)
Pred vs P2
CheXpert 0.958 (0.068) 42.1 (49.4) 24.3 (37.5) 0.966 (0.070) 21.5 (34.7) 16.9 (28.5)
ChestX-Ray8 0.969 (0.077) 33.3 (51.6) 16.6 (35.1) 0.965 (0.105) 18.2 (46.8) 16.2 (44.0)
MIMIC-CXR-JPG 0.962 (0.062) 36.9 (42.6) 18.3 (26.2) 0.945 (0.083) 30.1 (43.6) 26.2 (39.6)
Padchest 0.974 (0.070) 27.2 (38.1) 11.9 (22.6) 0.979 (0.037) 14.4 (26.4) 11.7 (21.9)
VinDr-CXR 0.982 (0.039) 22.5 (35.3) 8.0 (16.0) 0.957 (0.102) 21.1 (47.3) 19.3 (44.5)
P1 vs P2
CheXpert 0.962 (0.035) 43.9 (55.9) 23.0 (37.8) 0.918 (0.064) 40.4 (29.6) 33.5 (25.0)
ChestX-Ray8 0.976 (0.031) 27.9 (25.5) 12.2 (14.1) 0.942 (0.064) 24.5 (20.4) 21.2 (19.0)
MIMIC-CXR-JPG 0.966 (0.036) 29.0 (21.8) 15.8 (12.9) 0.932 (0.059) 31.5 (22.4) 26.6 (19.9)
Padchest 0.974 (0.027) 30.8 (23.1) 13.3 (12.5) 0.951 (0.038) 25.3 (19.8) 20.8 (16.3)
VinDr-CXR 0.983 (0.024) 22.7 (18.5) 8.1 (9.0) 0.943 (0.037) 25.3 (15.4) 21.6 (14.5)
@@
Figure 7: Histograms illustrating the distribution of segmentation quality across the five datasets. Each histogram represents the distribution of RCA (Reverse Classification Accuracy) estimations of DSC for a specific dataset, providing a visual representation of the segmentation performance. The histograms are truncated at 0.65 for visualization purposes. The full range of values is considered in the statistical analysis (Table 3).
@@
Figure 8 presents histograms that clearly demonstrate a superior RCA-estimated DSC for PA images compared to AP images across all investigated datasets. This is consistent with the fact that AP images tend to come from hospitalized patients, who are more difficult to position in standard views and usually include artifacts or cables that may act as confounders or occlude body parts, making anatomical segmentation more challenging. In contrast, discrepancies based on sex were less pronounced, with a slightly diminished segmentation quality observed in females. This is also consistent with the fact x-ray imaging of the upper thorax women’s breasts tend to occlude the imaged organs, resulting in poorer image contrast for the relevant anatomy [34] and making the segmentation task more challenging. Furthermore, the influence of disease findings on segmentation quality varied amongst datasets, with significant disparities evident in the ChestX-ray8 and CheXpert datasets.
@@
Table 3 can be useful to estimate the prevalence of such out-of-distribution images in each dataset. Notably, the VinDr-CXR dataset does not include any non-posteroanterior (PA) images nor many out-of-distribution images, as all images were manually annotated by physicians. However, the inclusion of zoomed-out images capturing both arms, which were absent in the HybridGNet training set, leads to lower-quality segmentations in those cases. In contrast, the larger datasets contain a higher number of cases mislabeled as PA or AP in their metadata. These mislabeled images could be easily identified and discarded by thresholding the RCA-estimated DSC, for example using the 1% or 5% quartile of a dataset as threshold. This approach can help to improve the overall quality and reliability of downstream applications developed using this datasets.
Figure 9: Illustration of a random subset of images with some of the lowest RCA values from each dataset. These images were identified as having low-quality segmentations based on their corresponding RCA scores. The selection showcases various types of images, including those with noise, mislabeled lateral CXR, and even non-medical objects like a mobile phone in the Padchest dataset. Reproduction of these images was allowed upon request to the original sources.
@@
We performed a final quality control at the dataset-level to evaluate the quality of HybridGNet segmentations per dataset (not individually as we did with RCA-estimated DSC). The hypothesis is that if the segmentations are good, they can be used as training set for a segmentation model that will obtain high performance. Thus, we trained five new HybridGNet models, by using the set of predicted masks of each of the five public datasets as training GT. We then evaluated model performance on the complete ChestXray-Landmarks database. All models were trained using the same configurations and hyperparameters as the original HybridGNet model [24]. As a comparison reference, we include the performance metrics reported on prior studies for the 20% test split of the ChestXray-Landmarks database which can be seen as an upper bound in performance, as this model was trained with the training split of the ChestXray-Landmarks database.
@@
The CheXmask dataset is intended to serve as a resource for researchers working in the field of medical imaging, particularly those interested in computational anatomy, shape understanding, CXR analysis and image segmentation. The dataset opens avenues for various practical applications, such as deep learning model development and evaluation, training of generative models, clinical decision support systems, disease detection and diagnosis, and anomaly detection. Researchers can explore novel methodologies, such as the integration of masked autoencoders for self-supervised learning, wherein the anatomical segmentation masks serve as a rich source for unsupervised feature extraction, or in the implementation of attention and deep supervision techniques guided by masks, facilitating enhanced model focus on specific anatomical regions for improved localization and feature extraction. As an example of clinical application, the task of cardiothoracic ratio estimation, can be particularly enhanced by leveraging the precise anatomical information encoded in the segmentation masks, thereby contributing to more robust and interpretable cardiomegaly detection methodologies.
For the sake of completeness, we release segmentation masks for all images included in our study. However, based on the qualitative and quantitative analyses presented in this paper (see section Technical Validation), our recommendation for downstream tasks is to only use segmentations whose RCA-estimated DSC is higher than 0.7, to avoid including out-of-distribution images as well as low quality masks. In addition, more attention should be put on specific subgroups (like pediatric patients in ChestX-ray8 or images acquired in AP position) for which we observed lower RCA-estimated DSC values overall, as discussed in Section "Validation via RCA-estimated DSC". Please also note that, when using segmentation masks for downstream tasks (e.g. computing cardiothoracic ratio from heart and lung segmentation masks), additional validation considering ground-truth annotations for the particular task should be performed to ensure the quality of the segmentation masks is good enough for the task at hand (e.g. manual measurements of cardiothoracic ratio should be employed to validate the quality of the segmentation masks).